OpenStudy (anonymous):

jake is nine years older than leth. three years ago, jake was twice as old as leth, how old are the two?

OpenStudy (anonymous):

y=x+9 y-3= 2(x-3) x=12 y=21

OpenStudy (mathmale):

Sounds like we're comparing the ages of others to Leth's age. Why not represent Leth's age by x? Then, Jake's age is x+9. Three years ago, Jake was twice as old as Leth. This can be translated into the following: Jake's age Leth's age x+9-3 x-3 Three years ago, Jake was twice as old as Leth. Thus, x+9-3 = 2 (x-3) The trick is to represent the unknowns properly. Again, start by choosing an appropriate variable to represent Leth's age. Can you now solve this for x? Then determine Leth's age and Jake's age.

OpenStudy (anonymous):

thank you :)

OpenStudy (mathmale):

My pleasure!

OpenStudy (anonymous):

:D